# Asymptotic safety, quantum gravity, and the swampland: a conceptual assessment

### v2

@@ 1 Introduction
Towards extending the swampland question beyond string theory — Within the swampland program, theories that become inconsistent when coupled to gravity are said to “belong to the swampland”, as opposed to the “landscape” of gravitational EFT which afford a consistent UV completion. Although the swampland program primarily originated and developed in the context of string theory, it sparked interest in the broader research community, and connections have been found in QG across approaches. Swampland criteria, such as the weak gravity conjecture Arkani-Hamed:2006emk ; Cheung:2018cwt ; Hamada:2018dde ; Aharony:2021mpc ; Harlow:2022ich ; Cordova:2022rer , the distance conjecture Ooguri:2006in ; Stout:2021ubb ; Stout:2022phm and the no-global-symmetries conjecture Misner:1957mt ; Banks:2010zn ; Harlow:2018tng ; Harlow:2018jwu ; McNamara:2019rup ; McNamara:2020uza ; McNamara:2021cuo ; McNamaraThesis ; Yonekura:2020ino ; Bah:2022uyz ; Bah:2024ucp constrain gravitational EFT with consistent UV completions, albeit unfortunately with a constraining power seemingly inversely proportional to the degree with which such criteria are established Palti:2019pca ; vanBeest:2021lhn ; Agmon:2022thq , see also section 2. As outlined above, whether such consistent theories are exclusively stringy is an open question, a positive answer to which is dubbed “string lamppost principle” Montero:2020icj ; Bedroya:2021fbu ; Bedroya:2023tch . Investigating these ideas beyond the context of string theory deAlwis:2019aud ; Basile:2021krr ; Knorr:2024yiu ; Eichhorn:2024rkc and comparing them to other top-down candidates would provide non-trivial tests of their validity, of string universality, or of the consistency of any given top-down candidate. In this paper, we make substantial steps in this research endeavor. We will restrict our discussion to QFT-based approaches, in particular asymptotically safe quantum gravity (ASQG), since they provide a direct connection to EFT at low energies. [^2] We focus on spacetime dimensions $d>3$. In $d\leq 3$ gravitons do not exist and gravity is topological, avoiding the arguments against a QFT description. In $d=3$, gravity in anti-de Sitter (AdS) sectors can admit a Chern-Simons Carlip:1995zj or, more precisely, a Virasoro Maloney:2007ud ; Collier:2023fwi description as a field theory. In $d=2$, Jackiw-Teitelboim gravity is UV complete but behaves as a statistical ensemble of theories Stanford:2019vob ; Kolchmeyer:2023gwa ; Penington:2023dql ; Penington:2024sum , and accordingly lacks black-hole microstates. Variations of this theory afford a non-unitary worldsheet description Collier:2023cyw ; Collier:2024kmo ; Collier:2024kwt ; Collier:2024lys ; Collier:2024mlg ; Collier:2025pbm ; Collier:2025lux . Our main focus will be relating ASQG to a subset of swampland ideas, accounting for their logical (co)implications and the arguments substantiating them. Nonetheless, the arguments presented in this paper, perhaps excluding those based on amplitudes as in section 4.1, apply to any QFT-based and Lorentz-invariant UV completion. We will mostly refer to string theory as a source of examples where such swampland ideas are realized, while motivating them without referring to any specific UV completion.
@@
A note on “effective” ASQG — When reflecting on the relationship between the string landscape and ASQG, it is noteworthy to highlight an alternative scenario to fundamental or “strict” ASQG, in which the fixed point marks the UV completion of the theory. Namely, one can conceive a scenario of “effective” asymptotic safety Percacci:2010af ; Eichhorn:2011pc ; Eichhorn:2019ybe ; deAlwis:2019aud ; Held:2020kze ; Eichhorn:2020sbo ; Basile:2021euh ; Basile:2021krk ; Eichhorn:2022ngh , in which an intermediate regime between the IR and a non-field-theoretic UV is controlled by some (possibly non-unitary) fixed point, close to which renormalization group (RG) trajectories can be approximated by ASQG. Such a scenario would evade all of our ensuing considerations, which strictly rest on requiring field-theoretic physics at all scales. Outside the context of this paper, investigating the possible intersections (if any) between the string landscape, effective asymptotic safety and observational constraints may lead to fruitful links between these fields, as well as phenomenology. This is illustrated in figure 1. For our purposes, we would like to emphasize that the potential issues for a field-theoretic description of gravity raised by swampland arguments are wholly avoided in effective ASQG. If this scenario were realized, one would be able to approximate e.g. Wilson coefficients of consistent EFT by those predicted by the RG flow stemming from a fixed point, any inconsistency of which would not impact the consistency of the full theory. [^6] As an example, Planck-scale suppressed violations of positivity bounds in ASQG Knorr:2024yiu , if confirmed, could be interpreted as hints of this scenario being realized. Such a setting would allow one to leverage the tools developed in the ASQG literature within an effective framework.
@@
Paper content — The contents of this paper are organized according to the overarching division between kinematic and dynamical considerations on QG. In section 2, we discuss some more detailed aspects of the swampland program and its consequences for EFT, together with some examples. In section 3, we begin by discussing kinematic aspects of the swampland. By this, we mean any consideration that does not involve solving equations of motion, diagonalizing Hamiltonians, and any such dynamical ingredients. Rather, the kinematic structure of QG (at least in the EFT regime) consists of spacetime and gauge-bundle topology, (generalized) symmetries, and the types of fields and charges. In this context, we discuss how completeness and the absence of global symmetries arise from spacetime topology fluctuations, holography, and black-hole entropy. We explain the logical connections between the various arguments and their consequences for ASQG. In section 4, we move on to dynamical aspects, which are less well-understood. We begin with the nature of observables in QG, their impact on weakly coupled UV completions and how ASQG can avoid it. Then we introduce the notions of infinite-distance limits and towers of species, their bottom-up motivations and the implications for field theories of gravity. As we shall see, such towers of species can only be consistent with ASQG in a particular case, and present some loopholes accordingly. Let us stress once more that our aim is to provide a general account of these ideas from a strictly bottom-up perspective, invoking string theory only to provide relevant examples. A middle-ground role in this sense is played by the anti-de Sitter/conformal field theory (AdS/CFT) correspondence, which can be viewed both in a top-down sense Maldacena:1997re and as a concrete general framework to study QG theories with stable, asymptotically AdS sectors, e.g. in the spirit of Montero:2017mdq ; Harlow:2018jwu ; Harlow:2018tng ; Montero:2018fns ; Baume:2020dqd ; Perlmutter:2020buo ; Baume:2023msm ; Perlmutter:2024noo ; Ooguri:2024ofs . We shall conclude in section 5 with a summary of our arguments, conclusions, and loopholes. We explain some of the more technical aspects and terminology in appendix A. As already remarked, to make the life of the reader easier, throughout the manuscript the words and concepts explained in the dictionary are highlighted and linked to the corresponding explanation in the dictionary.
@@ 2 The swampland program and asymptotic safety
The above results are based on anomaly cancellation. However, combining this condition with swampland conjectures arising from infinite-distance limits implies further bounds, e.g. that maximally supersymmetric $SU(N)$ gauge theory in four dimensions cannot be consistently coupled to (four-dimensional) gravity for $N>23$ Kim:2019ths . This bound is a consequence of the statistical properties of black objects and of the swampland distance conjecture, which we will review in section 4.2. The former is a more robust assumption but, as we shall discuss below, a consistent ASQG scenario would have to evade it. Thus, if such a loophole exists, the arguments of Kim:2019ths would not exclude consistent theories with $N>23$ in ASQG.
@@ 3 The kinematic swampland
By and large, the best understood and developed aspects of the swampland program encompass the kinematics of QG McNamaraThesis . The interplay between spacetime topology fluctuations Hawking:1978pog ; Gibbons:1991tp ; McNamara:2019rup ; McNamara:2021cuo ; McNamara:swamplandia , holography tHooft:1993dmi ; Susskind:1994vu , the absence of (generalized Gaiotto:2014kfa ; Schafer-Nameki:2023jdn ; Brennan:2023mmt ; Bhardwaj:2023kri ; Shao:2023gho ; Costa:2024wks ) symmetries Misner:1957mt ; Harlow:2018jwu ; Harlow:2018tng ; Heidenreich:2020pkc ; Heidenreich:2021xpr ; McNamara:2021cuo and completeness of gauge charge spectra Polchinski:2003bq ; Banks:2010zn ; Harlow:2021trr ; Kang:2022orq form a tightly woven web of theoretical constraints, depicted in figure 2. How this web interfaces with approaches to QG other than string theory is an underexplored question, and in this section we are going to analyze the compatibility between this web of kinematic swampland principles (cf. figure 2) and QG theories based on QFT in spacetime, such as ASQG. We will unveil a potential structural difference between the principles of the kinematic swampland conjectures and those of strictly field-theoretical descriptions QG; we will also clearly identify the root cause of such a difference. As we shall see, the consistency of strict ASQG (cf. figure 1) or other QFT-based approaches seems to require a non-thermodynamic interpretation of semiclassical black-hole physics. This conclusion is supported by various lines of reasoning based on topology change and the thermodynamic behavior of black holes and related ideas, which we are going to motivate and discuss next. These considerations also interweave with the dynamical considerations that we shall present in section 4.1. We will also comment on loopholes and generalizations of ASQG that could avoid this conclusion.
Figure 2: Various connections between the kinematic aspects of the swampland are shown. The dashed arrow denotes a weaker connection, indicating holography from the Bekenstein-Hawking area law. When specializing to AdS/CFT holography, one can prove Harlow:2018jwu ; Harlow:2018tng completeness and the absence of global symmetries. The blue region in the lower part of the diagram contains connections which persist without summing over spacetime topologies. In appendix A we provide explanations for some of the technical jargon.
@@ 3.1 Topology change in quantum gravity and field theory
The notion of summing over spacetime topologies, and whether it is required by consistency of QG, has been debated since its inception Hawking:1978pog ; Gibbons:1991tp . Topology change lies at the heart of the kinematical swampland program (cf. figure 2), and hence its realizability in QG plays a key role in unraveling the interface between (kinematic) swampland conjectures and ASQG. For this reason, in the following we begin our discussion by reviewing some of the arguments that have been proposed which support topology change in QG, emphasizing their strong connections with black-hole thermodynamics, the ideas of no global symmetries Misner:1957mt , holography, and the completeness of charges. The key question is thus whether (strict) ASQG and spacetime topology change are compatible with each other. Here we provide arguments to the effect that, assuming that ASQG is consistent, its spacetime topology is fixed. This does not forbid spatial topology change over time. Rather, by spacetime topology change, we specifically refer to summing over spacetime topologies in the path integral. As emphasized in the introduction, our work brings together a substantial amount of technical literature, mostly coming from the string community, with the ideas at the core of ASQG. It is beyond our scope to review all the underlying results, but we will attempt to describe the intuitive reasoning and provide references for more details.
@@
The third argument is more technical. The sum over topologies can be argued to reflect a sort of redundancy of QG Jafferis:2017tiu ; Marolf:2020xie ; McNamaraThesis ; McNamara:swamplandia , which can be understood in terms of the possible relation between gravity and entanglement in the sense of $ER=EPR$. In order for this to be possible, in the “gauge-fixed” description in which no explicit sum over topologies appears, the degrees of freedom of the theory must be able to replace its effects. A crucial ingredient in this respect is the triviality of cobordism classes McNamara:2019rup , which also follows from the holographic principle McNamaraThesis and subsumes the absence of (generalized) symmetries in gravity. [^12] To see this, on the one hand one can observe that non-trivial cobordism classes $[M]\in\Omega_{k}$ would provide conserved charges under a $(d-k-1)$-form symmetry Gaiotto:2014kfa group $\Omega_{k}^{\vee}$, which are physically carried by gravitational solitons McNamara:2021cuo kinematically described as connected sums $\mathbb{R}^{d-k-1,1}\times(\mathbb{R}^{k}\#M)$. On the other hand, such objects break (and abelianize) non-invertible symmetries, and provide the necessary representations to achieve gauge charge completeness McNamara:2021cuo . The absence of global symmetries can be independently motivated and related to completeness, as we shall explain in detail below. In summary, the third motivation for topology change is the $ER=EPR$ correspondence and the holographic principle.
@@
To summarize, topology change is motivated by an extensive network of arguments, but they seem very hard to implement in a QFT of gravity on spacetime, since quantum fields live, by definition, on spacetimes of fixed topology. In the absence of topology change, the network of swampland ideas in figure 2 is left with the blue-shaded region: black-hole thermodynamics, completeness of charges, and the absence of global symmetries — which we are going to discuss next — are still required by consistency based on semiclassical arguments.
@@ 3.3 Consequences of no topology change
In section 3.1 we argued that the strictly field-theoretic formulation of ASQG and spacetime topology change are hardly compatible. We shall now examine some consequences of this possibility, which more generally ought to apply to any QG theory without topology change. In the absence of topology change, the network of intertwined principles depicted in figure 2 reduces to the blue-shaded region. The absence of topology change thereby has consequences for the arguments leading to the absence of global symmetries, charge completeness, and black-hole entropy that we discuss subsequently.
@@
In the absence of spacetime topology change, the most robust arguments to deduce the absence of any and all global symmetries, [^18] Be it group-like, higher-form, categorical, non-invertible — perhaps the most general such notion is embodied by a (higher) fusion category of topological operators, up to the finiteness axiom for simple objects. completeness of gauge charge spectra Polchinski:2003bq ; Banks:2010zn and their equivalence Heidenreich:2020pkc ; Heidenreich:2021xpr ; McNamara:2021cuo due to gravitational solitons McNamara:2021cuo , and cobordism triviality McNamara:2019rup ; Ooguri:2020sua are lost. Dynamically, contributions to processes violating global charge conservation Bah:2022uyz and black-hole entropy Gibbons:1976ue ; Balasubramanian:2022gmo ; Geng:2024jmm due to non-trivial topologies are absent. Does this mean that global symmetries are allowed and gauge charge spectra may not be complete if ASQG is consistent? As we discussed in section 3.2, one can argue for either constraint purely on grounds of semiclassical black-hole thermodynamics Misner:1957mt , without invoking holography, topology change, or UV physics (except for unitarity). However, without the solid interconnected network of logical implications provided by topology fluctuations, see the blue-shaded region in figure 2, the resulting case is less compelling than in the full picture. However, as we shall see in the next point, black-hole thermodynamics needs a different interpretation, and thus the close relationship with global symmetries is lost.
@@
The potential presence of global symmetries in ASQG is supported by how RG flows interact with Ward identities when the regulator preserves a non-anomalous symmetry. More precisely, Ward identities show that no symmetry-breaking terms are generated by the flow if the regulator does not break the symmetry Gies:2006wv . This has been confirmed in various explicit computations for global symmetries of kinetic terms. For example, shift symmetry Narain:2009fy ; Percacci:2015wwa ; Eichhorn:2017als ; Laporte:2021kyp , chiral symmetry Eichhorn:2011pc ; Eichhorn:2016vvy ; Meibohm:2016mkp ; Eichhorn:2017eht ; deBrito:2020dta , or $O(N)$-symmetry Labus:2015ska ; deBrito:2021pyi ; Eichhorn:2021qet ; Eichhorn:2025ezh are not broken by gravitational fluctuations. This means that, first, all explored symmetry-breaking terms can be consistently set to zero at all scales, and second, that the symmetric subspace has an interacting UV fixed point, see also Eichhorn:2022gku ; Eichhorn:2025ezh for a discussion. [^19] This does not mean that, in general, all fixed points admit global symmetries, or that one could not break it dynamically even if one starts at a fixed point with a global symmetry. There is also a first piece of evidence from lattice computations that a global shift symmetry of quenched scalars is unbroken Dai:2021fqb . Another observation related to (the lack of) symmetries is that the pattern of generalized $p$-form symmetries extends down to $p=-1$, where a $(-1)$-form “symmetry” encodes a free parameter/coupling in the theory McNamara:2020uza . Since UV-critical surfaces typically have non-zero dimensions, it would be interesting to assess the role of these “symmetries” in ASQG.
@@
by virtue of the difficulty in reproducing an area law for a mixed state which is available in the EFT without local boundary dynamics (see also Buoninfante:2021ijy ), [^20] In consistent QG theories that include an asymptotically AdS sector, the connection between the area law and holographic dynamics is starkly visible via the AdS/CFT correspondence Maldacena:1997re ; Witten:1998qj , where (non-extremal) black holes are encoded in bona fide (thermal) states of a local conformal field theory. In this sense, the (grand-)canonical partition function computed within the EFT is, in fact, an actual thermodynamic quantity. the epistemically most economical route for ASQG seems to involve a significant alteration in how one should interpret black-hole thermodynamics. A popular but somewhat superficial argument Shomer:2007vq states that the presence of a UV fixed point implies that the density of high-energy states behaves like that of a conformal field theory in the bulk (thus scaling as a volume, rather than as an area). Potential loopholes of this argument Doboszewski:2017gim ; Bonanno:2020bil include large anomalous scalings, as well as the option of realizing scale-invariant QFT that are not conformally invariant El-Showk:2011xbs , although, in the relativistic setting, no interacting example has been found so far in four dimensions and above. Alternatively, one could try giving up IR locality (which might also be relevant in relation to avoiding global symmetries Borissova:2024hkc ), although it may bring along other issues, related to additional degrees of freedom, Wick rotation, and observations. The conclusion that black-hole thermodynamics needs a reinterpretation in UV-complete gravitational QFT is one of the main upshots of our analysis, and we will further corroborate it in section 4.

### v3

@@ 1 Introduction
Towards extending the swampland question beyond string theory — Within the swampland program, theories that become inconsistent when coupled to gravity are said to “belong to the swampland”, as opposed to the “landscape” of gravitational EFT which afford a consistent UV completion. Although the swampland program primarily originated and developed in the context of string theory, it sparked interest in the broader research community, and connections have been found in QG across approaches. Swampland criteria, such as the weak gravity conjecture Arkani-Hamed:2006emk ; Cheung:2018cwt ; Hamada:2018dde ; Aharony:2021mpc ; Harlow:2022ich ; Cordova:2022rer , the distance conjecture Ooguri:2006in ; Stout:2021ubb ; Stout:2022phm and the no-global-symmetries conjecture Misner:1957mt ; Banks:2010zn ; Harlow:2018tng ; Harlow:2018jwu ; McNamara:2019rup ; McNamara:2020uza ; McNamara:2021cuo ; McNamaraThesis ; Yonekura:2020ino ; Bah:2022uyz ; Bah:2024ucp constrain gravitational EFT with consistent UV completions, albeit unfortunately with a constraining power seemingly inversely proportional to the degree with which such criteria are established Palti:2019pca ; vanBeest:2021lhn ; Agmon:2022thq , see also section˜2. As outlined above, whether such consistent theories are exclusively stringy is an open question, a positive answer to which is dubbed “string lamppost principle” Montero:2020icj ; Bedroya:2021fbu ; Bedroya:2023tch . Investigating these ideas beyond the context of string theory deAlwis:2019aud ; Basile:2021krr ; Knorr:2024yiu ; Eichhorn:2024rkc and comparing them to other top-down candidates would provide non-trivial tests of their validity, of string universality, or of the consistency of any given top-down candidate. In this paper, we make substantial steps in this research endeavor. We will restrict our discussion to QFT-based approaches, in particular asymptotically safe quantum gravity (ASQG), since they provide a direct connection to EFT at low energies. [^2] We focus on spacetime dimensions $d>3$. In $d\leq 3$ gravitons do not exist and gravity is topological, avoiding the arguments against a QFT description. In $d=3$, gravity in anti-de Sitter (AdS) sectors can admit a Chern-Simons Carlip:1995zj or, more precisely, a Virasoro Maloney:2007ud ; Collier:2023fwi description as a field theory. In $d=2$, Jackiw-Teitelboim gravity is UV complete but behaves as a statistical ensemble of theories Stanford:2019vob ; Kolchmeyer:2023gwa ; Penington:2023dql ; Penington:2024sum , and accordingly lacks black-hole microstates. Variations of this theory afford a non-unitary worldsheet description Collier:2023cyw ; Collier:2024kmo ; Collier:2024kwt ; Collier:2024lys ; Collier:2024mlg ; Collier:2025pbm ; Collier:2025lux . Our main focus will be relating ASQG to a subset of swampland ideas, accounting for their logical (co)implications and the arguments substantiating them. Nonetheless, the arguments presented in this paper, perhaps excluding those based on amplitudes as in section˜4.1, apply to any QFT-based and Lorentz-invariant UV completion. We will mostly refer to string theory as a source of examples where such swampland ideas are realized, while motivating them without referring to any specific UV completion.
@@
A note on “effective” ASQG — When reflecting on the relationship between the string landscape and ASQG, it is noteworthy to highlight an alternative scenario to fundamental or “strict” ASQG, in which the fixed point marks the UV completion of the theory. Namely, one can conceive a scenario of “effective” asymptotic safety Percacci:2010af ; Eichhorn:2011pc ; Eichhorn:2019ybe ; deAlwis:2019aud ; Held:2020kze ; Eichhorn:2020sbo ; Basile:2021euh ; Basile:2021krk ; Eichhorn:2022ngh , in which an intermediate regime between the IR and a non-field-theoretic UV is controlled by some (possibly non-unitary) fixed point, close to which renormalization group (RG) trajectories can be approximated by ASQG. Such a scenario would evade all of our ensuing considerations, which strictly rest on requiring field-theoretic physics at all scales. Outside the context of this paper, investigating the possible intersections (if any) between the string landscape, effective asymptotic safety and observational constraints may lead to fruitful links between these fields, as well as phenomenology. This is illustrated in figure˜1. For our purposes, we would like to emphasize that the potential issues for a field-theoretic description of gravity raised by swampland arguments are wholly avoided in effective ASQG. If this scenario were realized, one would be able to approximate e.g. Wilson coefficients of consistent EFT by those predicted by the RG flow stemming from a fixed point, any inconsistency of which would not impact the consistency of the full theory. [^6] As an example, Planck-scale suppressed violations of positivity bounds in ASQG Knorr:2024yiu , if confirmed, could be interpreted as hints of this scenario being realized. Such a setting would allow one to leverage the tools developed in the ASQG literature within an effective framework.
@@
Paper content — The contents of this paper are organized according to the overarching division between kinematic and dynamical considerations on QG. In section˜2, we discuss some more detailed aspects of the swampland program and its consequences for EFT, together with some examples. In section˜3, we begin by discussing kinematic aspects of the swampland. By this, we mean any consideration that does not involve solving equations of motion, diagonalizing Hamiltonians, and any such dynamical ingredients. Rather, the kinematic structure of QG (at least in the EFT regime) consists of spacetime and gauge-bundle topology, (generalized) symmetries, and the types of fields and charges. In this context, we discuss how completeness and the absence of global symmetries arise from spacetime topology fluctuations, holography, and black-hole entropy. We explain the logical connections between the various arguments and their consequences for ASQG. In section˜4, we move on to dynamical aspects, which are less well-understood. We begin with the nature of observables in QG, their impact on weakly coupled UV completions and how ASQG can avoid it. Then we introduce the notions of infinite-distance limits and towers of species, their bottom-up motivations and the implications for field theories of gravity. As we shall see, such towers of species can only be consistent with ASQG in a particular case, and present some loopholes accordingly. Let us stress once more that our aim is to provide a general account of these ideas from a strictly bottom-up perspective, invoking string theory only to provide relevant examples. A middle-ground role in this sense is played by the anti-de Sitter/conformal field theory (AdS/CFT) correspondence, which can be viewed both in a top-down sense Maldacena:1997re and as a concrete general framework to study QG theories with stable, asymptotically AdS sectors, e.g. in the spirit of Montero:2017mdq ; Harlow:2018jwu ; Harlow:2018tng ; Montero:2018fns ; Baume:2020dqd ; Perlmutter:2020buo ; Baume:2023msm ; Perlmutter:2024noo ; Ooguri:2024ofs . We shall conclude in section˜5 with a summary of our arguments, conclusions, and loopholes. We explain some of the more technical aspects and terminology in appendix˜A. As already remarked, to make the life of the reader easier, throughout the manuscript the words and concepts explained in the dictionary are highlighted and linked to the corresponding explanation in the dictionary.
@@ 2 The swampland program and asymptotic safety
The above results are based on anomaly cancellation. However, combining this condition with swampland conjectures arising from infinite-distance limits implies further bounds, e.g. that maximally supersymmetric $SU(N)$ gauge theory in four dimensions cannot be consistently coupled to (four-dimensional) gravity for $N>23$ Kim:2019ths . This bound is a consequence of the statistical properties of black objects and of the swampland distance conjecture, which we will review in section˜4.2. The former is a more robust assumption but, as we shall discuss below, a consistent ASQG scenario would have to evade it. Thus, if such a loophole exists, the arguments of Kim:2019ths would not exclude consistent theories with $N>23$ in ASQG.
@@ 3 The kinematic swampland
By and large, the best understood and developed aspects of the swampland program encompass the kinematics of QG McNamaraThesis . The interplay between spacetime topology fluctuations Hawking:1978pog ; Gibbons:1991tp ; McNamara:2019rup ; McNamara:2021cuo ; McNamara:swamplandia , holography tHooft:1993dmi ; Susskind:1994vu , the absence of (generalized Gaiotto:2014kfa ; Schafer-Nameki:2023jdn ; Brennan:2023mmt ; Bhardwaj:2023kri ; Shao:2023gho ; Costa:2024wks ) symmetries Misner:1957mt ; Harlow:2018jwu ; Harlow:2018tng ; Heidenreich:2020pkc ; Heidenreich:2021xpr ; McNamara:2021cuo and completeness of gauge charge spectra Polchinski:2003bq ; Banks:2010zn ; Harlow:2021trr ; Kang:2022orq form a tightly woven web of theoretical constraints, depicted in figure˜2. How this web interfaces with approaches to QG other than string theory is an underexplored question, and in this section we are going to analyze the compatibility between this web of kinematic swampland principles (cf. figure˜2) and QG theories based on QFT in spacetime, such as ASQG. We will unveil a potential structural difference between the principles of the kinematic swampland conjectures and those of strictly field-theoretical descriptions QG; we will also clearly identify the root cause of such a difference. As we shall see, the consistency of strict ASQG (cf. figure˜1) or other QFT-based approaches seems to require a non-thermodynamic interpretation of semiclassical black-hole physics. This conclusion is supported by various lines of reasoning based on topology change and the thermodynamic behavior of black holes and related ideas, which we are going to motivate and discuss next. These considerations also interweave with the dynamical considerations that we shall present in section˜4.1. We will also comment on loopholes and generalizations of ASQG that could avoid this conclusion.
Figure 2: Various connections between the kinematic aspects of the swampland are shown. The dashed arrow denotes a weaker connection, indicating holography from the Bekenstein-Hawking area law. When specializing to AdS/CFT holography, one can prove Harlow:2018jwu ; Harlow:2018tng completeness and the absence of global symmetries. The blue region in the lower part of the diagram contains connections which persist without summing over spacetime topologies. In appendix˜A we provide explanations for some of the technical jargon.
@@ 3.1 Topology change in quantum gravity and field theory
The notion of summing over spacetime topologies, and whether it is required by consistency of QG, has been debated since its inception Hawking:1978pog ; Gibbons:1991tp . Topology change lies at the heart of the kinematical swampland program (cf. figure˜2), and hence its realizability in QG plays a key role in unraveling the interface between (kinematic) swampland conjectures and ASQG. For this reason, in the following we begin our discussion by reviewing some of the arguments that have been proposed which support topology change in QG, emphasizing their strong connections with black-hole thermodynamics, the ideas of no global symmetries Misner:1957mt , holography, and the completeness of charges. The key question is thus whether (strict) ASQG and spacetime topology change are compatible with each other. Here we provide arguments to the effect that, assuming that ASQG is consistent, its spacetime topology is fixed. This does not forbid spatial topology change over time. Rather, by spacetime topology change, we specifically refer to summing over spacetime topologies in the path integral. As emphasized in the introduction, our work brings together a substantial amount of technical literature, mostly coming from the string community, with the ideas at the core of ASQG. It is beyond our scope to review all the underlying results, but we will attempt to describe the intuitive reasoning and provide references for more details.
@@
The third argument is more technical. The sum over topologies can be argued to reflect a sort of redundancy of QG Jafferis:2017tiu ; Marolf:2020xie ; McNamaraThesis ; McNamara:swamplandia , which can be understood in terms of the possible relation between gravity and entanglement in the sense of $ER=EPR$. In order for this to be possible, in the “gauge-fixed” description in which no explicit sum over topologies appears, the degrees of freedom of the theory must be able to replace its effects. A crucial ingredient in this respect is the triviality of cobordism classes McNamara:2019rup , which also follows from the holographic principle McNamaraThesis and subsumes the absence of (generalized) symmetries in gravity. [^12] To see this, on the one hand one can observe that non-trivial cobordism classes $M]\in\Omega_{k}$ would provide conserved charges under a (dk1)(d-k-1)-form symmetry Gaiotto:2014kfa group $\Omega_{k}^{\vee}$, which are physically carried by gravitational solitons McNamara:2021cuo kinematically described as connected sums $\mathbb{R}^{d-k-1,1}\times(\mathbb{R}^{k}\#M)$. On the other hand, such objects break (and abelianize) non-invertible symmetries, and provide the necessary representations to achieve gauge charge completeness McNamara:2021cuo . The absence of global symmetries can be independently motivated and related to completeness, as we shall explain in detail below. In summary, the third motivation for topology change is the $ER=EPR$ correspondence and the holographic principle.
@@
To summarize, topology change is motivated by an extensive network of arguments, but they seem very hard to implement in a QFT of gravity on spacetime, since quantum fields live, by definition, on spacetimes of fixed topology.  [^16] In lattice simulations of ASQG, in the form of dynamical triangulations, the topology is fixed explicitly, as the performed Pachner moves are by construction topology-preserving PACHNER1991129 . In the absence of topology change, the network of swampland ideas in figure˜2 is left with the blue-shaded region: black-hole thermodynamics, completeness of charges, and the absence of global symmetries — which we are going to discuss next — are still required by consistency based on semiclassical arguments.
@@ 3.3 Consequences of no topology change
In section˜3.1 we argued that the strictly field-theoretic formulation of ASQG and spacetime topology change are hardly compatible. We shall now examine some consequences of this possibility, which more generally ought to apply to any QG theory without topology change. In the absence of topology change, the network of intertwined principles depicted in figure˜2 reduces to the blue-shaded region. The absence of topology change thereby has consequences for the arguments leading to the absence of global symmetries, charge completeness, and black-hole entropy that we discuss subsequently.
@@
In the absence of spacetime topology change, the most robust arguments to deduce the absence of any and all global symmetries, [^19] Be it group-like, higher-form, categorical, non-invertible — perhaps the most general such notion is embodied by a (higher) fusion category of topological operators, up to the finiteness axiom for simple objects. completeness of gauge charge spectra Polchinski:2003bq ; Banks:2010zn and their equivalence Heidenreich:2020pkc ; Heidenreich:2021xpr ; McNamara:2021cuo due to gravitational solitons McNamara:2021cuo , and cobordism triviality McNamara:2019rup ; Ooguri:2020sua are lost. Dynamically, contributions to processes violating global charge conservation Bah:2022uyz and black-hole entropy Gibbons:1976ue ; Balasubramanian:2022gmo ; Geng:2024jmm due to non-trivial topologies are absent. Does this mean that global symmetries are allowed and gauge charge spectra may not be complete if ASQG is consistent? As we discussed in section˜3.2, one can argue for either constraint purely on grounds of semiclassical black-hole thermodynamics Misner:1957mt , without invoking holography, topology change, or UV physics (except for unitarity). However, without the solid interconnected network of logical implications provided by topology fluctuations, see the blue-shaded region in figure˜2, the resulting case is less compelling than in the full picture. However, as we shall see in the next point, black-hole thermodynamics needs a different interpretation, and thus the close relationship with global symmetries is lost.
@@
The potential presence of global symmetries in ASQG is supported by how RG flows interact with Ward identities when the regulator preserves a non-anomalous symmetry. More precisely, Ward identities show that no symmetry-breaking terms are generated by the flow if the regulator does not break the symmetry Gies:2006wv . This has been confirmed in various explicit computations for global symmetries of kinetic terms. For example, shift symmetry Narain:2009fy ; Percacci:2015wwa ; Eichhorn:2017als ; Laporte:2021kyp , chiral symmetry Eichhorn:2011pc ; Eichhorn:2016vvy ; Meibohm:2016mkp ; Eichhorn:2017eht ; deBrito:2020dta , or $O(N)$-symmetry Labus:2015ska ; deBrito:2021pyi ; Eichhorn:2021qet ; Eichhorn:2025ezh are not broken by gravitational fluctuations. This means that, first, all explored symmetry-breaking terms can be consistently set to zero at all scales, and second, that the symmetric subspace has an interacting UV fixed point, see also Eichhorn:2022gku ; Eichhorn:2025ezh for a discussion. [^20] This does not mean that, in general, all fixed points admit global symmetries, or that one could not break it dynamically even if one starts at a fixed point with a global symmetry. There is also a first piece of evidence from lattice computations that a global shift symmetry of quenched scalars is unbroken Dai:2021fqb . Another observation related to (the lack of) symmetries is that the pattern of generalized $p$-form symmetries extends down to p=1p=-1, where a (1)(-1)-form “symmetry” encodes a free parameter/coupling in the theory McNamara:2020uza . Since UV-critical surfaces typically have non-zero dimensions, it would be interesting to assess the role of these “symmetries” in ASQG.
@@
by virtue of the difficulty in reproducing an area law for a mixed state which is available in the EFT without local boundary dynamics (see also Buoninfante:2021ijy ), [^21] In consistent QG theories that include an asymptotically AdS sector, the connection between the area law and holographic dynamics is starkly visible via the AdS/CFT correspondence Maldacena:1997re ; Witten:1998qj , where (non-extremal) black holes are encoded in bona fide (thermal) states of a local conformal field theory. In this sense, the (grand-)canonical partition function computed within the EFT is, in fact, an actual thermodynamic quantity. the epistemically most economical route for ASQG seems to involve a significant alteration in how one should interpret black-hole thermodynamics. A popular but somewhat superficial argument Shomer:2007vq states that the presence of a UV fixed point implies that the density of high-energy states behaves like that of a conformal field theory in the bulk (thus scaling as a volume, rather than as an area). Potential loopholes of this argument Doboszewski:2017gim ; Bonanno:2020bil include large anomalous scalings, as well as the option of realizing scale-invariant QFT that are not conformally invariant El-Showk:2011xbs , although, in the relativistic setting, no interacting example has been found so far in four dimensions and above. Alternatively, one could try giving up IR locality (which might also be relevant in relation to avoiding global symmetries Borissova:2024hkc ), although it may bring along other issues, related to additional degrees of freedom, Wick rotation, and observations. The conclusion that black-hole thermodynamics needs a reinterpretation in UV-complete gravitational QFT is one of the main upshots of our analysis, and we will further corroborate it in section˜4.